OpenStudy (boxman61):

Need help with process, im stumped x+6/2 + x+2/3 =9

OpenStudy (boxman61):

\[\frac{ x+6 }{ 2 }+\frac{ x+2 }{ 3 }=9\] Answer = 32/5 Unknown how to reach the correct answer

Nnesha (nnesha):

first find common denominator

OpenStudy (boxman61):

6?

Nnesha (nnesha):

yes right \[\huge\rm \frac{ 3(x+6)+2(x+2) }{ 6 }=9\] multiply the numerator of first fraction by the denominator of 2nd fraction multiply the numerator of 2nd fraction by the denominator of 1st fraction 1st) multiply both sides by 6 2nd) distribute parentheses by outside number

OpenStudy (jhannybean):

\[\frac{ x+6 }{ 2 }+\frac{ x+2 }{ 3 }=\frac{9}{\color{red}{1}}\]This might help, too!

OpenStudy (boxman61):

\[\frac{ 5x+22 }{ 6 }=9\] \[6\times9=54\] \[54-22=32\] \[32\div5=\frac{ 32 }{ 5 }\]

Nnesha (nnesha):

great job!!!!

OpenStudy (boxman61):

i realize that was probably not the proper way but thats my life story... thank you both

OpenStudy (jhannybean):

As long as you understand the method, everything works. I was a little confused at first since you hadnt written out your steps but after solving it,it looks good :)